How many calories are required to heat each of the following from 15 °C to 65 °C? (a) 3.0g of aluminum, (b) 5.0g of pyrex glass, (c) 20g of platinum. The specific heats, in cal/g C, for aluminum, pyrex, and platinum are 0.21, 0.20, and 0.032, repectively.


1
Expert's answer
2022-03-31T11:49:17-0400
"Q=cm\Delta t"

(a)

"Q=0.21*3*50=32\:\rm J"

(b)

"Q=0.20*5*50=50\:\rm J"

(b)

"Q=0.032*20*50=32\:\rm J"